# MotionLab: Unified Human Motion Generation and Editing via the Motion-Condition-Motion Paradigm

### v4

@@ Appendix A Additional Quantitative Results
CondMDI (Cohan et al., 2024) 1 0.1551 0.6787 9.5807 0.0936 0.3739
@@
Table 7. Evaluation of motion in-between with CondMDI (Cohan et al., 2024) on HumanML3D (Guo et al., 2022a) dataset.
Also as shown in Figure 8, our framework also outperform MCM-LDM on all metrics, demonstrating the effectiveness of our framework in motion style transfer.
Figure 8. Comparison of the motion style transfer with MCM-LDM (Song et al., 2024) on a subset of HumanML3D (Guo et al., 2022a). This shows that our model has a stronger ability to preserve the semantics of source motion and a stronger ability to learn the style of style motion.
Appendix B Additional Ablation Studies
@@
To further validate the Aligned ROPE, we also introduce the variant of 3D-Learnable and 3D-ROPE to distinguish the source motion, target motion and trajectory. As shown in Table 8 and Figure 9, 1D-position encoding are better than 3D-position encoding by avoiding introducing distances between different modalities, and ROPE are better than learnable position encoding by explicit positional encoding. Hence our 1D-ROPE outperforms all others variants, demonstrating its effective to embed the position information into tokens.
To further validate the motion curriculum learning, we adopt the variant of removing the masked pre-training and directly supervised fine-tuning in order; the variant of with masked pre-training but supervised fine-tuning all tasks together; the variant of introducing masked reconstruction, motion in-between and trajectory based motion generation in orderly. As shown in Table 9, our proposed motion curriculum learning outperforms all other variants, highlighting the effective of masked pre-training and fine-tuning tasks in order by avoiding gradient conflicts between different tasks. Specifically, the variant of masked pre-training in order demonstrates that necessity of introduce motion in-between and trajectory-based motion generation together, or will greatly weakens the performance of the model in the latter task.
@@
Figure 9. Ablation results of MotionLab on the motion in-between (with text). Beige motion is use 1D-learnable position encoding, purple motion use Aligned ROPE, and gray motions are the poses provided in keyframes, demonstrating the importance of Aligned ROPE.
@@
Table 8. Ablation studies of our MotionLab’s position encoding on each task.
@@
Table 9. Ablation studies of our MotionLab’s motion curriculum learning on each task.
@@
Table 10. Ablation studies of our MotionLab’s task combinations.
Appendix C Representation for Each Modality
We represent the features of all modalities as tokens for the attention mechanism (Vaswani, 2017). Specifically, source motion and target motion are represented as $M_{S}\in\mathbb{R}^{N\times D}$ and $M_{T}\in\mathbb{R}^{N\times D}$, and we first ignore timestep $t$ here. For the instruction, it is represented as $I\in\mathbb{R}^{1\times 768}$ extracted from the CLIP (Radford et al., 2021). For available conditions $C$, the text is represent as $p\in\mathbb{R}^{77\times 768}$ extracted from the last hidden layer of CLIP, the trajectory is represented as $h\in\mathbb{R}^{N\times J\times 3}$, and the style is represented as $s\in\mathbb{R}^{1\times 512}$ extracted from Zhong et al..
Appendix D Instrcutions for Each Task
As shown in the Table13, the instructions in the Task Instruction Modulations for each task are presented, which benefits our framwork to distinguish differents tasks.
@@ Appendix D Instrcutions for Each Task
Table 11. Instructions in the Task Instruction Modulations for each task.
Appendix E Classifier Free Guidance for Each Task
Classifier-Free Guidance (CFG) (Ho and Salimans, 2022) has been incorporated for various tasks (Peng et al., 2024; Zhang et al., 2025b, 2024) based on diffusion models. As shown in Table12, strengths of classifier free guidance for each task are presented, which contributing to the results’ quality during sampling.
@@ Appendix E Classifier Free Guidance for Each Task
Table 12. Strength of classifier free guidance for each task.
Appendix F 3D Assets
Task Instruction
in-between (without text) “generate motion by given key frames.”
style-based generation “generate motion by given style.”
trajectory-based editing “edit source motion by given trajectory.”
text-based editing “edit source motion by given text.”
style transfer “generate motion by the given style and content.”
text-based generation “generate motion by given text.”
Table 13. Instructions in the Task Instruction Modulations for each task.

### v5

@@ 8 Details of Rectified Flows
$x_{t-\frac{1}{N}}=x_{t}-\frac{1}{N}v_{\theta}(t,x_{t})$ (7)
where $N$ is the discretization number of the interval [0,1].
9 MotionLab Inference
During inference, Classifier-Free Guidance (CFG) [27] is incorporated for both motion generation and motion editing to boost sampling quality and align conditions and target motion.
$v_{\theta}(M_{T},t,C)=$ $v_{\theta}(M_{T}|t,\emptyset)+$
$\lambda_{C}[v_{\theta}(M_{T}|t,C)-v_{\theta}(M_{T}|t,\emptyset)]$ (8)
where $t$ is the timestep and $\lambda_{C}>1$ is a hyper-parameter to control the strength of the corresponding conditional guidance.
$v_{\theta}(M_{T},t,M_{S},C)=$ $v_{\theta}(M_{T}|t,\emptyset,\emptyset)$
$+\lambda_{S}[v_{\theta}(M_{T}|t,S,\emptyset)-v_{\theta}(M_{T}|t,\emptyset,\emptyset)]$
$+\lambda_{C}[v_{\theta}(M_{T}|t,S,C)-v_{\theta}(M_{T}|t,S,\emptyset)]$ (9)
10 Memory Usage and Time Cost
The maximum memory usage during training is 23 GB for each GPU. The memory usage and the time spent during inference are summarized in the following Table 7.
Metric text gen traj. gen text edit traj. edit in-between style transfer
memory usage (GB) 4.16 4.31 5.83 6.81 4.32 5.74
time spend (AITS) 0.068 0.134 0.160 0.191 0.142 0.152
Table 7: The memory usage and time cost of MotionLab.
11 Additional Quantitative Results
As shown in Table 8, our framework outperforms CondMDI on all settings, illustrating the effectiveness of our framework in motion in-between.
[…]
CondMDI [11] 1 0.1551 0.6787 9.5807 0.0936 0.3739
@@ 11 Additional Quantitative Results
Table 8: Evaluation of motion in-between with CondMDI [11] on HumanML3D [21] dataset.
Also, as shown in Figure 8, our framework also outperforms MCM-LDM on all metrics, demonstrating the effectiveness of our framework in motion style transfer.
Figure 8: Comparison of the motion style transfer with MCM-LDM [55] on a subset of HumanML3D [21]. This shows that our model has a stronger ability to preserve the semantics of source motion and a stronger ability to learn the style of style motion.
12 Additional Ablation Studies
@@
To further validate the Aligned ROPE, we also introduce the variant of 3D-Learnable and 3D-ROPE to distinguish the source motion, target motion, and trajectory. As shown in Table 10 and Figure 9, 1D-position encoding is better than 3D-position encoding by avoiding introducing distances between different modalities, and ROPE are better than learnable position encoding by explicit positional encoding. Hence, our 1D-ROPE outperforms all other variants, demonstrating its effectiveness in embedding the position information into tokens.
To further validate the motion curriculum learning, we adopt three variants: removing the masked pre-training and directly supervised fine-tuning in order; with masked pre-training but supervised fine-tuning all tasks together; and introducing masked reconstruction, motion in-between, and trajectory-based motion generation in an orderly manner. As shown in Table 11, motion curriculum learning outperforms all other variants, highlighting the effectiveness of masked pre-training and fine-tuning tasks in order to avoid gradient conflicts between different tasks. Specifically, the variant of masked pre-training in order to demonstrate the necessity of introducing motion in-between and trajectory-based motion generation together, or it will greatly weaken the performance of the model in the latter task.
The explanation of “w/o task instruction modulation” uses ’null’ as the instruction for all tasks, rather than learned task tokens or one-hot encoding vectors. We have conducted an additional ablation experiment to examine these situations, which can be suboptimal due to the random initialization of their parameters, as shown in Table 9.
one-hot encoding 0.187 0.0369 56.18 71.52 0.0287 43.20 66.98
learnable tokens 0.183 0.0356 56.03 71.89 0.0288 41.20 64.98
Ours 0.167 0.0334 56.34 72.65 0.0283 44.62 69.21
Table 9: Ablation studies of Task Instruction Modulation.
@@
Figure 9: Ablation results of MotionLab on the motion in-between (with text). Beige motion is use 1D-learnable position encoding, purple motion use Aligned ROPE, and gray motions are the poses provided in keyframes, demonstrating the importance of Aligned ROPE.
@@
Table 10: Ablation studies of our MotionLab’s position encoding on each task.
@@
Table 11: Ablation studies of our MotionLab’s motion curriculum learning on each task.
@@
Table 12: Ablation studies of our MotionLab’s task combinations.
13 Representation for Each Modality
We represent the features of all modalities as tokens for the attention mechanism [60]. Specifically, source motion and target motion are represented as $M_{S}\in\mathbb{R}^{N\times D}$ and $M_{T}\in\mathbb{R}^{N\times D}$, and we first ignore timestep $t$ here. For the instruction, it is represented as $I\in\mathbb{R}^{1\times 768}$ extracted from the CLIP [49]. For available conditions $C$, the text is represent as $p\in\mathbb{R}^{77\times 768}$ extracted from the last hidden layer of CLIP, the trajectory is represented as $h\in\mathbb{R}^{N\times J\times 3}$, and the style is represented as $s\in\mathbb{R}^{1\times 512}$ extracted from [71].
14 Instructions for Each Task
@@
Table 13: Instructions in the Task Instruction Modulations for each task.
As shown in the Table 13, the instructions in the Task Instruction Modulations for each task are presented, which benefits our framework to distinguish different tasks.
15 Classifier Free Guidance for Each Task
@@
Table 14: Strength of classifier free guidance for each task.
As shown in Table 14, strengths of classifier-free guidance for each task are presented, which contribute to the results’ quality during sampling. We conduct ablation experiments based on the hyperparameters provided by the baseline and finally obtain the above hyperparameters.
16 3D Assets